OpenStudy (anonymous):

joes java makes a blend of coffee which is uses two coffees priced at $8.40 /pound & $7.28/pounds perspectively. how much of the first should be used to mix with 9 pounds of the second, if the mixture will sell at $7.95/pound? write the equation that best represents?

OpenStudy (anonymous):

thanks yuhiluvv yuh

OpenStudy (anonymous):

ithink that too cos iono

OpenStudy (aripotta):

here, try this: 8.4x + 7.28y = 7.95(x + 9)

OpenStudy (aripotta):

ugh. idk. i'm sorry

OpenStudy (aripotta):

8.4x + 7.28(9) = 7.95(x + 9)

OpenStudy (anonymous):

A. 840x + 7.28(9 - x) = 7.95(9) B. 8.40x + 7.28(9) = 7.95(x+9) C. 8.40(9) + 7.28x = 7.95 answer chioces

OpenStudy (aripotta):

aha! there it is!

OpenStudy (aripotta):

b

OpenStudy (anonymous):

ahaha thanks love ur tha best!!!

OpenStudy (aripotta):

;)